OpenStudy (anonymous):

sin theta =-4/5 cos beta=-5/13 Find the exact value of tan(theta+beta) Will FAN and MEDAL I know the formula I need help with the steps

OpenStudy (anonymous):

@zepdrix do you know how to do this?

zepdrix (zepdrix):

Remember that when you apply your Tangent Angle Addition Formula, it's going to give you a bunch of `tan(theta)` and `tan(beta)` showing up. So* what we're doing right now is drawing a triangle, and trying to figure out what `tan(theta)` is.

OpenStudy (anonymous):

For Tan theta I have -4/3 and for tan beta I have -5/12

OpenStudy (anonymous):

5^2 instead of 15 from the given

zepdrix (zepdrix):

Oh why do I have a 15? 0_o oh boy...

zepdrix (zepdrix):

So from this:\[\Large\rm \color{royalblue}{\sin \theta=-\frac{4}{5}},\qquad \color{orangered}{\cos \beta=-\frac{5}{13}}\]You determined,\[\Large\rm \color{royalblue}{\tan \theta=-\frac{4}{3}},\qquad \color{orangered}{\tan \beta=-\frac{12}{5}}\]I think maybe the tan beta is off, lemme check really quick.

zepdrix (zepdrix):

you determined -5/12, but i think it's -12/5* is what i meant to say.

zepdrix (zepdrix):